# Controllable Learning in Information Retrieval: A Survey

### v1

@@ Abstract
Controllable learning (CL) emerges as a critical component in trustworthy machine learning, ensuring that learners meet predefined targets and can adaptively adjust without retraining according to the changes in those targets. We provide a formal definition of CL, and discuss its applications in information retrieval (IR) where information needs are often complex and dynamic. The survey categorizes CL according to who controls (users or platforms), what is controllable (e.g., retrieval objectives, users’ historical behaviors, controllable environmental adaptation), how control is implemented (e.g., rule-based method, Pareto optimization, Hypernetwork), and where to implement control (e.g.,pre-processing, in-processing, post-processing methods). Then, we identify challenges faced by CL across training, evaluation, task setting, and deployment in online environments. Additionally, we outline promising directions for CL in theoretical analysis, efficient computation, empowering large language models, application scenarios and evaluation frameworks in IR.
Index Terms:
Controllable learning, trustworthy machine learning, information retrieval
@@ 1 Introduction
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=base east, child anchor=base west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=13em,fill=fill_0,draw=draw_0,font=,, where level=1text width=9.5em,fill=fill_1,draw=draw_1,font=,, where level=2text width=11.4em,fill=fill_2,draw=draw_2,font=,, where level=3text width=12.4em,fill=fill_3,draw=draw_3,font=,, where level=4font=,, [ Controllable Learning (CL)
@@
(Sec. 3) [ Who controls
@@
(Sec. 3.1.1) [ Prior profile modification which empowers users to actively shape their recommendation experience: UCRS [1], IFRQE [2], LACE [3], and  [4]. , leaf, text width=31.8em ] ] [ Platform-mediated control
(Sec. 3.1.2) [ Impose algorithmic adjustments and policy-based constraints on the recommendation process: ComiRec [5], CMR [6], CCDF [7] and  [8]. , leaf, text width=31.8em ] ] ] [ What is controllable
(Sec. 3.2) [ Multi-Objective Control
(Sec. 3.2.1) [ Taking several objectives (metrics) as control targets, these work includes ComiRec [5], UCRS [1], CMR [6] , leaf, text width=31.8em ] ] [ Historical Behavior Control
(Sec. 3.2.2) [ Taking historical behavior as control targets , these work includes LACE [3], LFRQE [2], Grapevine [4] , leaf, text width=31.8em ] ] [ Controllable Environmental
Adaptation(Sec. 3.2.3) [ Taking environment adaptation as target, these work includes HyperBndit[9] and so on. , leaf, text width=31.8em ] ] ] ] [ How control is
@@
(Sec. 4.1.3) [ Hypernetworks based methods: Hamur[24], CMR[6], Hyperprompt[25], HyperBandit[9], [26], [27], [16], and [17] , leaf, text width=31.8em ] ] ] [ Where to control
@@
(Sec. 4.2.2) [ Including regularization and constrained optimization methods:CMR [6], CCDF [7]. , leaf, text width=31.8em ] ] [ Post-processing methods
(Sec. 4.2.3) [ Including reranking and result diversification methods: ComiRec [5] and UCRS [1]. , leaf, text width=31.8em ] ] ] ] [ Evaluations for CL
@@
Trustworthiness in machine learning is becoming increasingly important, with considerations such as controllability, fairness, privacy, and interpretability gaining prominence [38, 39, 40]. In this paper, we focus on the controllability of machine learning methods and their applications. Norbert Wiener, the originator of cybernetics, foresaw the ethical challenge of learning machine over sixty years ago when he stated [41]As machines learn they may develop unforeseen strategies at rates that baffle their programmers.” Additionally, Wiener explained the significance of controllability in learning machines [41]: “If we use, to achieve our purposes, a mechanical agency with whose operation we cannot efficiently interfere once we have started it, because the action is so fast and irrevocable that we have not the data to intervene before the actions complete, then we had better be quite sure that the purpose put into the machine is the purpose which we really desire and not merely a colorful imitation of it, etc.”, emphasizing two fundamental aspects of AI controllability: (1) A controllable learning machine needs to ensure that its predictions/decisions meet the target of the AI user; (2) Intervening in a learning machine after its deployment is a crucial but challenging task.
@@
Controllability has been extensively discussed in the field of generative machine learning models, also known as controllable generation. For instance, in text generation [42, 43, 44] or visual generation [45, 46, 47], the AI models take a given task description (often referred to as a “prompt”) as input and generates content that aligns with that description. However, in the realm of discriminative machine learning models and applications, there lacks a unified definition and in-depth discussion of controllability. The current popular prompt-based controllable generation methods can be seen as a form of pre-processing, where, under fixed model parameters, changing input features allows the model to meet the target of the AI user. Existing theoretical results suggest that controlling model parameters may lead to better convergence than controlling model inputs when given task requirements [27], revealing that achieving controllability in machine learning involves various methods beyond prompt-based approaches. Therefore, controllable learning (CL) is a crucial category worthy of comprehensive review and exploration in trustworthy machine learning, ensuring that learning models meet user expectations and adhere to ethical standards. Meanwhile, in the context of Information Retrieval (IR) applications, CL enables models to adapt to diverse task description (e.g, users’ information needs) at test time without extensive retraining, delivering personalized and relevant results. Given the above landscape, this survey attempts to provide a formal definition of controllable learning, classify and discuss its application paradigms in information retrieval. In summary, this survey focuses on the topic of controllable learning, which aims to find a learner capable of adapting to different task requirements without the need for retraining, thereby meeting the desired task targets of the AI users.
@@
In the context of MaaS, the demand for controllability of ML models in Information Retrieval (IR) applications becomes more pronounced. This is attributed to the inherent complexity of information needs in IR tasks, where precise and tailored results are critical. In traditional settings, such as keyword-based search engines, the retrieval process is relatively straightforward, relying on predefined rules and algorithms. However, with the advent of MaaS and the utilization of large language models, the landscape has evolved. Now, users expect finer-grained and context-aware responses based on their information needs. This heightened expectation emphasizes the importance of controllability of ML models. Specifically, there is a need for mechanisms that enable individuals or platforms to easily express and convey their specific requirements for IR tasks. For example, in an e-commerce setting, a user searching for “black T-shirt” may expect different results based on varying weights of evaluation metrics (such as favoring diversity to simultaneously expose clothing that pairs well with black T-shirts), or setting specific filters to exclude irrelevant content (such as not exposing previously purchased T-shirts). Achieving this level of controllability necessitates sophisticated ML models capable of dynamically adjusting model parameters or outputs based on task requirements, without extensive retraining. Therefore, in the realm of IR, MaaS not only provides unprecedented access to powerful ML models but also underscores the significance of controllability and customization in addressing the diverse needs of users across different domains and applications.
@@ 2 Foundations of Controllable Learning
The concept of controllable learning is manifested across various AI research directions, yet there is currently lack of a unified formal definition. To explicitly express the ingredients of controllable learning and unify existing controllable learning methods, we first define controllable learning in terms of task specifications. In brief, controllable learning is the ability to find a learner that can adapt to different task requirements without the need for retraining, thereby meeting the desired task targets, as detailed in Definition 1.
Definition 1 (Controllable Learning).
@@
The integration of the learner $f$ and the control function $h$ is called a controllable learner. Moreover, upon receiving a new task requirement $\mathcal{T}^{\prime}\in\Gamma$, the control function $h$ should be capable of outputting a new learner $f_{\mathcal{T}^{\prime}}$ without the need for model training, ensuring that $f_{\mathcal{T}^{\prime}}$ satisfies the task target $\bm{s}^{\prime}_{\mathrm{tgt}}\in\mathcal{T}^{\prime}$.
@@
For the task requirement in Definition 1, the task target $\bm{s}_{\mathrm{tgt}}$ can be considered as the ideal quantitative metric that the controllable learner aims to achieve. The task description $\bm{s}_{\mathrm{desc}}$ is the specific representation of $\bm{s}_{\mathrm{tgt}}$ that can be inputted into an algorithm. $\bm{s}_{\mathrm{desc}}$ can be represented as a natural language prompt or as a vector containing multiple weights, among other forms. The context $\bm{s}_{\mathrm{ctx}}$ in task requirements can consist of features like historical data and the current user’s profile, offering controllable learners more background knowledge for prediction.
Under the definition, the procedure of controllable learning could be shown as Figure 2. When the output space $\mathcal{Y}$ in Definition 1 is not labels or lists but instead generated content, it is referred to as controllable generation. In this survey, our primary focus lies on controllable learning methods within information retrieval applications, which typically belong to controllable discrimination rather than controllable generation.
Figure 2: The procedure of controllable learning in Definition 1.
@@ 3 Controllable Learning in Information Retrieval: Taxonomy
Controllable learning plays a crucial role in modern information retrieval (IR), where learning models must dynamically adapt to various task descriptions, such as users’ specific information needs, without requiring extensive retraining. This capability ensures the delivery of personalized and relevant search results, thereby enhancing user satisfaction in IR systems. Recent advancements underscore the importance of controllability in contemporary IR platforms [8, 1, 3]. The following sections delve into the taxonomy of controllable learning in IR, exploring who controls and what is controllable in IR systems. Additionally, we also present some representative works and provide a detailed description of their control methods in terms of “who” and “what” categories. (Table I).

### v2

# A Survey of Controllable Learning: Methods and Applications in Information Retrieval

@@ Abstract
Controllability has become a crucial aspect of trustworthy machine learning, enabling learners to meet predefined targets and adapt dynamically at test time without requiring retraining as the targets shift. We provide a formal definition of controllable learning (CL), and discuss its applications in information retrieval (IR) where information needs are often complex and dynamic. The survey categorizes CL according to what is controllable (e.g., multiple objectives, user portrait, scenario adaptation), who controls (users or platforms), how control is implemented (e.g., rule-based method, Pareto optimization, hypernetwork and others), and where to implement control (e.g., pre-processing, in-processing, post-processing methods). Then, we identify challenges faced by CL across training, evaluation, task setting, and deployment in online environments. Additionally, we outline promising directions for CL in theoretical analysis, efficient computation, empowering large language models, application scenarios and evaluation frameworks.
keywords:
controllable learning, information retrieval, model adaptation
[…]
@@ 1 Introduction
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=base east, child anchor=base west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=13em,fill=fill_0,draw=draw_0,font=,, where level=1text width=9.5em,fill=fill_1,draw=draw_1,font=,, where level=2text width=10.4em,fill=fill_2,draw=draw_2,font=,, where level=3text width=10.4em,fill=fill_3,draw=draw_3,font=,, where level=4font=,, [ Controllable Learning (CL)
@@
(Sec. 3) [ What is controllable
(Sec. 3.1) [ Multi-Objective Control
(Sec. 3.1.1) [ Taking several objectives (metrics) as control targets, these work includes ComiRec [1], UCRS [2], CMR [3], PadiRec [4], RecLM-gen [5] , leaf, text width=31.8em ] ] [ User Portrait Control
(Sec. 3.1.2) [ Taking historical behavior as control targets , these work includes LACE [6], LFRQE [7], Grapevine [8], TEARS [9], CMBR [10], UCR [11] , leaf, text width=31.8em ] ] [ Scenario Adaptation Control
(Sec. 3.1.3) [ Taking environment adaptation as target, these work includes HyperBndit[12], Hamur [13], HyperBandit [12], PEPNet [14] , leaf, text width=31.8em ] ] ] [ Who controls
@@
(Sec. 3.2.1) [ Prior profile modification which empowers users to actively shape their recommendation experience: UCRS [2], IFRQE [7], LACE [6], Supervised $\beta$-VAE [15], LP [16], TEARS [9], LangPTune [17],UCRS [2] , leaf, text width=31.8em ] ] [ Platform-mediated control
(Sec. 3.2.2) [ Impose algorithmic adjustments and policy-based constraints on the recommendation process: ComiRec [1], CMR [3], CCDF [18] and  [19], Padirec [4], SAMD [20], HyperBandit [12], DTRN [21] , leaf, text width=31.8em ] ] ] ] [ How control is
@@
(Sec. 4.1.3) [ Hypernetworks based methods: Hamur [13], CMR [3], Hyperprompt [36], HyperBandit [12], PadiRec [4], [37], [28], and [29] , leaf, text width=31.8em ] ] ] [ Where to control
@@
(Sec. 4.2.2) [ Including regularization and constrained optimization methods:CMR [3], CCDF [18], HyperBandit [12] , leaf, text width=31.8em ] ] [ Post-processing methods
(Sec. 4.2.3) [ Including reranking and result diversification methods: ComiRec [1] and MMR [39] , leaf, text width=31.8em ] ] ] ] [ Evaluations for CL
@@
Trustworthiness in machine learning is becoming increasingly important, with considerations such as controllability, fairness, privacy, and interpretability gaining prominence [49, 50, 51]. In November 2023, the Bletchley Declaration [52] , adopted at the first Global Artificial Intelligence Safety Summit, emphasized that:“AI should be designed, developed, deployed, and used, in a manner that is safe, in such a way as to be human-centric, trustworthy and responsible…Substantial risks may arise from potential intentional misuse or unintended issues of control relating to alignment with human intent.” In October 2023, the Cyberspace Administration of China released the Global AI Governance Initiative [53], which emphasized the need to: “…make AI technologies more secure, reliable, controllable, and equitable, …ensure that AI always remains under human control.” In this paper, we explore the topic of controllable AI, with a particular focus on the controllability of machine learning methods and their practical applications. Norbert Wiener, the originator of cybernetics, foresaw the ethical challenge of learning machine over sixty years ago when he stated [54]As machines learn they may develop unforeseen strategies at rates that baffle their programmers.” Additionally, Wiener explained the significance of controllability in learning machines [54]: “If we use, to achieve our purposes, a mechanical agency with whose operation we cannot efficiently interfere once we have started it, because the action is so fast and irrevocable that we have not the data to intervene before the actions complete, then we had better be quite sure that the purpose put into the machine is the purpose which we really desire and not merely a colorful imitation of it, etc.”, emphasizing two fundamental aspects of AI controllability: (1) A controllable learning machine needs to ensure that its predictions/decisions meet the target of the AI user; (2) Intervening in a learning machine after its deployment is a crucial but challenging task.
@@
Controllability has been extensively discussed in the field of generative machine learning models, also known as controllable generation. For instance, in text generation [55, 56, 57] or visual generation [58, 59, 60], the AI models take a given task description (often referred to as a “prompt”) as input and generates content that aligns with that description. However, in the realm of discriminative machine learning models and applications, there lacks a unified definition and in-depth discussion of controllability. The current popular prompt-based controllable generation methods can be seen as a form of pre-processing, where, under fixed model parameters, changing input features allows the model to meet the target of the AI user. Existing theoretical results suggest that controlling model parameters may lead to better convergence than controlling model inputs when given task requirements [37], revealing that achieving controllability in machine learning involves various methods beyond prompt-based approaches. Therefore, controllable learning (CL) is a crucial category worthy of comprehensive review and exploration in trustworthy machine learning, ensuring that learning models meet user expectations and adhere to ethical standards. Meanwhile, in the context of learning-based Information Retrieval (IR) applications, users actively search for information or passively receive recommendations, both aimed at fulfilling their information-seeking needs [61, 62]. In these applications, controllable learning (CL) enables models to adapt to to diverse task description (e.g, users’ information needs) at test time without extensive retraining, delivering personalized and relevant results. Given the above landscape, this survey attempts to provide a formal definition of controllable learning, classify and discuss its application paradigms in information retrieval. In summary, this survey focuses on the topic of controllable learning, which aims to find a learner capable of adapting to different task requirements without the need for retraining, thereby meeting the desired task targets of the AI users.
@@
In the context of MaaS, the demand for controllability of ML models in information retrieval applications becomes more pronounced. This is attributed to the inherent complexity of information needs in IR tasks, where precise and tailored results are critical. In traditional settings, such as keyword-based search engines, the retrieval process is relatively straightforward, relying on predefined rules and algorithms. However, with the advent of MaaS and the utilization of large language models, the landscape has evolved. Now, users expect finer-grained and context-aware responses based on their information needs. This heightened expectation emphasizes the importance of controllability of ML models. Specifically, there is a need for mechanisms that enable individuals or platforms to easily express and convey their specific requirements for IR tasks. For example, in an e-commerce setting, a user searching for “black T-shirt” may expect different results based on varying weights of evaluation metrics (such as favoring diversity to simultaneously expose clothing that pairs well with black T-shirts), or setting specific filters to exclude irrelevant content (such as not exposing previously purchased T-shirts). Achieving this level of controllability necessitates sophisticated ML models capable of dynamically adjusting model parameters or outputs based on task requirements, without extensive retraining. Therefore, in the realm of IR, MaaS not only provides unprecedented access to powerful ML models but also underscores the significance of controllability and customization in addressing the diverse needs of users across different domains and applications.
@@ 2 Foundations of Controllable Learning
The concept of controllable learning is manifested across various AI research directions, yet there is currently lack of a unified formal definition. To explicitly express the ingredients of controllable learning and unify existing controllable learning methods, we first define controllable learning in terms of task specifications. In brief, controllable learning is the ability to find a learner that can adapt to different task requirements without the need for retraining, thereby meeting the desired task targets, as detailed in Definition 2.1.
Definition 2.1 (Controllable Learning).
@@
The integration of the learner $f$ and the control function $h$ is called a controllable learner. Moreover, upon receiving a new task requirement $\mathcal{T}^{\prime}\in\Gamma$ at test time , the control function $h$ should be capable of outputting a new learner $f_{\mathcal{T}^{\prime}}$ without the need for model retraining, ensuring that $f_{\mathcal{T}^{\prime}}$ satisfies the task target $\bm{s}^{\prime}_{\mathrm{tgt}}\in\mathcal{T}^{\prime}$.
@@
For the task requirement in Definition 2.1, the task target $\bm{s}_{\mathrm{tgt}}$ can be considered as the ideal quantitative metric that the controllable learner aims to achieve. The task description $\bm{s}_{\mathrm{desc}}$ is the specific representation of $\bm{s}_{\mathrm{tgt}}$ that can be inputted into an algorithm. $\bm{s}_{\mathrm{desc}}$ can be represented as a natural language prompt or as a vector containing multiple weights, among other forms. The context $\bm{s}_{\mathrm{ctx}}$ in task requirements can consist of features like historical data and the current user’s profile, offering controllable learners more background knowledge for prediction.
Under the definition, the procedure of controllable learning could be shown as Figure 2. When the output space $\mathcal{Y}$ in Definition 2.1 is not labels or lists but instead generated content, it is referred to as controllable generation. In this survey, our primary focus lies on controllable learning methods within information retrieval applications, which typically belong to controllable discrimination rather than controllable generation.
Figure 2: The procedure of controllable learning in Definition 2.1. Specifically, the context contains the historical interaction data, user profile, and other side information. The task target can be divided into two categories based on the source: user side and platform side, which include factors such as user mood, user privacy, and the platform’s multi-objective requirements. The task description is a representation of the task target, such as the preference weights of multi objective, natural language and so on. The controllable learner consists of the control function $h$ and the leaner $f$, taking the task description and context as input and outputting a new learner $f_{\mathcal{T}}$ to accomplish recommendation consistent of the task target.
@@ 3 Controllable Learning in Information Retrieval: Taxonomy
Controllable learning plays a crucial role in modern information retrieval (IR), where learning models must dynamically adapt to various task descriptions, such as users’ specific information needs, without requiring extensive retraining. This capability ensures the delivery of personalized and relevant search results, thereby enhancing user satisfaction in IR systems. Recent advancements underscore the importance of controllability in contemporary IR platforms [19, 2, 6]. The following sections delve into the taxonomy of controllable learning in IR, exploring who controls and what is controllable in IR systems. Additionally, we also present some representative works and provide a detailed description of their control methods in terms of “what” and “who” categories. (Table 1).
3.1 What is Controllable
This section aims to clarify “what is controllable” for a controllable learner in information retrieval applications. Specifically, we explore the question: “What does the task target $\bm{s}_{\mathrm{tgt}}$ in Definition 2.1 of controllable learning could be?” The first type of task target is multi-objective control. In real-world business scenarios, both users and platforms experience shifts in interests, especially in multi-objective settings, where the emphasis on different objectives varies over time. Therefore, enabling the model to adapt to these changing objective requirements is particularly important. The second task target is user portrait control, where users can edit the context, thereby influencing the recommendation output. Some platforms provide preference questionnaires, allowing users to express personal preferences independently. Similarly, some platforms allow users to adjust their profiles and interaction history, achieving specific needs such as privacy, personalization and so on. The third task target is scenario adaptation control. Real-world platforms typically contain multiple content pages, where users’ transition data across different pages can be valuable. Additionally, in the time dimension, user behavior on some platforms is often specifically related to time segments (considered as a special scenario). Explicitly leveraging scenario information to control the recommendation system is therefore essential.
Following this classification, Section 3.1.1 provides a detailed analysis of existing work on multi-objective control. Section 3.1.2 focuses on user portrait control, and Section 3.1.3 summarizes existing research on scenario adaptation control.
3.1.1 Multi-Objective Control
The preference (task targets) of the platform for multiple objectives (such as accuracy, diversity, and novelty) change over time. When facing a new task target, we expect a control function to directly map the learner $f$ to $f_{\mathcal{T}}$ that meets the demand based on the task description $\bm{s}_{\mathrm{{desc}}}$, where the $\bm{s}_{\mathrm{{desc}}}$ is defined as the preference for each objectives (in the explicit form of weights or natural language or the implicit form of interaction behaviors and so on). Moreover, at a finer granularity, user preferences for different categories can also be regarded as a form of multi-objective control. An example is shown in Figure 3.
Figure 3: The work flow of the Multi-Objective Control. Over time, the objectives of both the platform and users change, leading to the emergence of multi-objective requirements.
ComiRec [1] strikes a balance between accuracy and diversity through hyperparameters in the aggregation module by capturing multiple user interests and integrating them into the recommendation process. UCRS [2] introduces a system where users can proactively reduce filter bubbles, controlling the recommendations they receive via hyperparameters that affect accuracy and coverage. The CMR framework [3] uses a hypernetwork to dynamically adjust preference weights, aiming to balance accuracy, diversity, etc. through platform-defined preference weights. These systems collectively focus on achieving a harmonious balance among the various objectives of recommendation quality, ensuring that users receive accurate, diverse, and relevant recommendations. PadiRec [4] treats multi-objective models under different preference weights as learning targets, considering the preference weights as conditions. Through conditional learning with a diffusion model, it captures the relationship between preference weights and the parameters of the multi-objective model. During the testing stage, by inputting the desired preference weights, the diffusion model can customize the model in real time according to these specifications, enabling control and test-time adaptation. RecLM-gen [5] utilize LLMs to follow recommendation-specific instructions. They utilize conventional recommendation models to distill recommendation capabilities into LLMs via labels. They then propose a reinforcement learning-based alignment procedure to enhance the generalization ability of LLMs, which includes category control and category proportion control.
3.1.2 User Portrait Control
On some platforms, the user profile or user history data can be edited to comply with privacy rules and provide users with greater control. Following our definition, this means that the context $\bm{s}_{\mathrm{ctx}}$ is editable, i.e., controllable, while in this case, the control function $h$ is a projection mapping that does not alter the learner $f$, specifically, $f=h(f,\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}})$. The task target is achieved by editing the input of the learner $f$. Some examples on the User Portrait Control are shown in Figure 4.
Figure 4: The examples on User Portrait Control. The user portrait is reflected in various aspects such as interaction history, preference summary, and ratings. The user portrait can be modified and edited. The deletions are marked in red, and additions are marked in green.
Some work in recommender systems involves allowing users to manage the context including their past interactions and profile information. LACE [6] allows users to edit their profiles by selecting or deselecting concepts, directly influencing the recommendations. IFRQE [7] lets users delete or modify past interactions to refine their recommendation experience. TEARS [9] summarizes user preferences in a text format, employing optimal transport to align summary-based embeddings with history-based black-box embeddings during the training stage. This allows users to edit their summaries, enabling them to control the recommendations they receive during the testing stage. Together, these methods empower users to shape their digital footprint and ensure that recommendations better align with their evolving interests and preferences. CMBR [10] is designed for the specific scenario of skin recommendations in games, leveraging multi-domain user information and various types of user behavior data. During the training phase, CMBR inserts a special token before each type of behavior in the sequence, enabling the sequential model to learn the relationship between each special token and its corresponding behavior. In the test phase, by appending the desired special token at the end of the sequence, the model can predict the specific behavior associated with that token among the multiple behavior types. UCR [11] proposes a user controllable recommendation framework with two essential properties. The first is retrospective controllability with retrospective explanation, which enables users to identify how their previous behaviors, such as clicks, purchases, and likes, contribute to current recommendations. The second is prospective controllability with prospective explanation, which offers users insight into how their interaction with the currently recommended items will influence future recommendations. Thus users can effectively control their future recommendations by operating in an informed manner.
Figure 5: The work flow of the Scenario Adaptation Control. As continuous scenario switching, the control function maps the learner to a scenario-specific learner based on the task description and the scenario-specific learner is then applied to the new scenario.
3.1.3 Scenario Adaptation Control
In real-world recommendation systems, platforms often consist of multiple content pages, thereby creating multiple scenarios. From a temporal perspective, different time segments can also be viewed as distinct scenarios. A unified model helps fully leverage information across these scenarios, but it typically requires control to adapt the model to the current scenario. Following our definition, $\bm{s}_{\mathrm{desc}}$ needs to include scenario-specific side information. Thus, we can use the control function $h$ to map the learner $f$ to a scenario-specific learner $f_{\mathcal{T}}$. The work flow of the Scenario Adaptation Control is shown in Figure 5. It is worth noting that some studies focus on transitions between multiple tasks, which we also consider as part of scenario adaptation.
Compared to dynamic online learning [72, 73, 74], controllable scenario adaptation is an explicit method that involves describing scenario-specific side information and inputting them into the controllable learner, rather than solely relying on the feedback of specific scenarios. Hamur [13] proposed a shared hypernetwork for dynamically generating the parameters of adapters to capture implicit information across different domains. The input to the hypernetwork consists of instance-level embeddings that carry domain-specific information, from which it generates adapter weight parameters. HyperBandit [12] utilizes a hypernetwork to explicitly model the binary relationship between periodic external environments and user interests. In the testing stage, HyperBandit takes time features as input and dynamically adjusts the recommendation model parameters to account for time-varying user preferences. PEPNet [14] introduces a gating mechanism that allows prior information to be injected into the model. This mechanism processes prior information with different personalized semantics through a two-layer neural network structure, generating personalized gating scores that adaptively control the importance of the prior information. DTRN [21] employing a hypernetwork to dynamically control task-behavior parameter interactions and refining representations with task-specific context awareness, DTRN effectively mitigates task interference and enhances representation specificity.